Assignment 12 Let A= -3 -5 -2 6 8 2 0 0 -2 [3pt] (a) Find the eigenvalues of A. (1pt] (b) Determine the algebraic multiplicithi, can you write the notation and everything?
Answer #1

0 су O -58 1-2 2 -2 0 o - 2 2 - 2nd 4 - O -2 2 multiplicity of each eigen value. A = 36 we have to find the eigen value of th- 5% + 6 y =D 0 - 21+ 24-42=0 54-by. x = 6 5. using in ® Haby -2168) +27-4220 942=-124+24 347= -29. 2=-_y 10 2) Ez = { 1 Gy;E-z={ 10,0,2) ; 261R}. basis or eigen space Wirilo 2202 spanf-2-{10,0,1); } is 19,0,1). dim Ez = 21A+2 f) 21 >> neometric muSo p= [ E3, E2, E-2 P 1 615 6 0 oto تا o 0 D = 0 O 2 Ar. o 0-2 Any doubt then then comment please. g will help you help you A